a garden plot is to contain 200 square feet. If its length is to be twice its width, what should its dimensions be?
Answer by mananth(16949)

Area = 200 ft^2
..
width be x
length = 2x
..
Area = 2x*x
2x^2= 200
/2
x^2=100
take square root
x= 10 the with.
length = 2x = 20 feet
...
20 feet by 10 feet
...
m.ananth@hotmail.ca